Grade 11Mechanics

The position of a particle is given by x=1+t-(t^2). The distance covered by the particle in 2 secs will be.

Khushi
8 Years agoGrade 11

1 Answer

swarup kumar majee
8 Years ago
As per the question x=1+t-(t2)
in two second x=1+2-(4)=1+2-4=-1
so the distance covered  is 1 in opposite direction